# Edge Zeta Functions and Eigenvalues for Spherical Buildings: A Uniform Hecke Algebra Approach

### v2

@@ 3.4 Luo’s decomposition theorem
The following subsection concerns a recent result of Luo [16, §2] on the decomposition of the longest element $w_{S}$ of a Weyl group. Luo originally formulated the theorem in the context of his work on the Casselman–Shahidi conjecture, where it serves as a key combinatorial input for analyzing the singularity of intertwining operators in exceptional types. In our setting the same theorem provides the structural simplification needed to collapse long products of relative destination elements to just a few double coset operators involving opposite parabolics.
@@
In terms of our notations, Luo’s original statement ([16]) is
@@
with length additivity. This is equivalent to $w_{S}w_{1}=(w^{\prime}_{1}w_{2})\cdots(w^{\prime}_{m-1}w_{m})$, which is in turn equivalent (by taking inverses and multiplying $(w^{\prime}_{0}w_{1})$ on the left) to the form
@@
used above. We adopt this form because it aligns more directly with the orbit structure of the geodesic edge graph and with our RDE computations. The values of $m$ for each irreducible type of $\Delta$ are listed in next section.
@@ 4 Examples: Explicit Zeta Factors for Finite Buildings
This section illustrates the general spectral and zeta–factor formulas developed in Section 3.2 by working out concrete cases for spherical buildings of finite classical groups. We focus on the split families of type $\mathbf{A}$ and type $\mathbf{C}$, for which all Hecke parameters equal $q$ and the zeta factors admit closed expressions. The type $\mathbf{A}$ case recalls the results of [18] in the present framework, while the type $\mathbf{C}$ case provides new explicit formulas for $G=\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$.
@@ 4.1 Finite buildings: split and non–split cases
By Tits’ classification of buildings, every finite spherical building of irreducible type and rank at least three arises from an algebraic group of Lie type over a finite field [21, §11.4]. A complete list of such families is given in [21, §11.2]. Our discussion applies uniformly to all of them.
@@
For twisted or non–split groups, the parameters $Q(s)$ vary between conjugacy classes of simple reflections. An important infinite family is provided by the unitary groups $^{2}A_{n}(q)=\mathrm{PGU}_{n+1}(q)$ ($n\geq 2$), defined by a non–degenerate hermitian form. Their building is the flag complex of the associated polar space, of type $\mathbf{C}_{\lfloor(n+1)/2\rfloor}$ [21, §11.2.2].
@@
$Z\!\bigl(X_{2}(B),u\bigr)=\prod_{C}Z_{C}(u),\qquad Z_{C}(u)=\prod_{\lambda\in\mathop{\rm Spec}\nolimits(T_{C})}\frac{1}{1-\lambda\,u^{\,c}}.$
@@
$Z\!\left(X_{2}(\mathcal{B})\!\mid_{C},u\right)=\prod_{\chi:\,n_{\chi}\neq 0}\;\prod_{\zeta^{d}=1}\frac{1}{\bigl(1-\zeta\,(Q(w_{I})^{-2}q^{\,f_{\chi}})^{1/d}\,u^{c}\bigr)^{m_{C,\chi}(\zeta)\,d_{\chi}}}.$
@@ 4.2 Type $\mathbf{A}$: recalling the formulas
We summarize the zeta functions for buildings of type $\mathbf{A}_{n}$ from [18] using the present three–step method; we refer to that paper for the details. By Tits’ classification [21, §11.2], all buildings of type $\mathbf{A}_{n}$ with $n\geq 3$ are associated with $G=\mathop{\rm GL}\nolimits_{n+1}(\mathbb{F}_{q})$.
@@
Step 3. Root selection. The eigenvalues of $T_{C}$ are the $(2m/c)$-th roots of the collapsed values above. In type $A$, $2m=6$.
@@
  • If $c=2$ (balanced), then $2m/c=3$: we need to take cubic roots of the eigenvalues $q^{\,f_{\lambda}-2\,\mathrm{wt}_{r}(\mu)}$. Multiplicities are as in Table 8 of [18, §8.3]: the $K_{\lambda,\mu}$ eigenvalues split as evenly as possible among $1,\omega,\omega^{2}$ ($\omega^{3}=1$). Thus
@@
    and $m(1)$ differs from $m(\omega)$ by at most $1$, according to $K_{\lambda,\mu}\bmod 3$.
Therefore we recover the following closed formulas in type $A$ [18, §7.2, §8.4].
@@
$Z_{C}\bigl(X_{2}(B),u\bigr)\;=\;\prod_{\lambda\vdash n}\frac{1}{\bigl(1-q^{\,f_{\lambda}-2\,\mathrm{wt}_{r}(\mu)}\,u^{6}\bigr)^{d_{\lambda}K_{\lambda,\mu}}}.$
@@
$\frac{1}{Z_{C}\!\bigl(X_{2}(B),u\bigr)}\;=\;\prod_{\lambda\vdash n}\;\prod_{j=0}^{2}\Bigl(1-\omega^{j}\,q^{(f_{\lambda}-2\,\mathrm{wt}_{r}(\mu))/3}\,u^{2}\Bigr)^{d_{\lambda}\,(K_{\lambda,\mu}+\epsilon_{\lambda,j})/3},$
@@ 4.3.1 Geometry and zeta factors set-up
$Z\!\bigl(X_{2}(\mathcal{B})\!\mid_{C},u\bigr)=\prod_{\chi:\,n_{\chi}\neq 0}\ \prod_{\zeta^{\,d}=1}\frac{1}{\Bigl(1-\zeta\,(Q(w_{I})^{-2}q^{\,f_{\chi}})^{1/d}\,u^{\,c}\Bigr)^{m_{C,\chi}(\zeta)\,d_{\chi}}}.$
@@
Beyond the split symplectic case $G=\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$, the same spherical buildings of type $\mathbf{C}$ arise from, among others, two other classical families [21, §11.2].
@@
Along a fixed type–orbit $C$ in the geodesic edge graph, the invariants $\ell(w_{I}),\,f_{\chi},\,n_{\chi},$ and $m_{C,\chi}(\zeta)$ depend only on the Weyl group $W(C_{n})$ and on the chosen orbit $C$; only the degree $d_{\chi}$ depends on the ambient group representation. For the natural correspondence between labels $(\lambda,\mu)\vdash n$ on $W(C_{n})$ and the representations of $G=\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$ or $\mathrm{O}_{2n+1}(\mathbb{F}_{q})$, the associated representations for $\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$ and $\mathrm{O}_{2n+1}(\mathbb{F}_{q})$ have the same dimensions $d_{\chi}$ [2, §13.8]. Consequently, the edge–geodesic zeta factors for $\mathcal{B}(\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q}))$ and $\mathcal{B}(\mathrm{O}_{2n+1}(\mathbb{F}_{q}))$ coincide, although the two buildings are not isomorphic when $n\geq 3$.
@@ 4.3.2 Spectral inputs for type $\mathbf{C}$
Let $S^{\lambda},S^{\mu}$ denote the Specht modules of the symmetric group $S_{n_{+}},S_{n_{-}}$, so its character is $\chi_{\lambda,1},\chi_{\mu,1}$ in §4.2. Inflate $S^{\lambda}\boxtimes S^{\mu}$ to $C_{2}^{\,n}\rtimes(S_{n_{+}}\times S_{n_{-}})$ (trivial on $C_{2}^{\,n}$) and set
@@
    $\mathop{\rm Irr}\nolimits(\mathcal{H}_{q})$ vs. $\mathop{\rm Irr}\nolimits(W_{n})$ The family $\{\mathcal{H}_{q}\}_{q\neq 0}$ is a flat deformation of $\mathbb{C}W_{n}$: for $q=1$ one has $\mathcal{H}_{1}=\mathbb{C}W_{n}$. By Tits’ deformation theorem [6, §8.1.7], there is a correspondence between irreducible representations
@@
    If $M$ is a simple $\mathbb{C}G$–module contained in $\mathbb{C}Ge_{B}$, then $e_{B}M$ is a simple $\mathcal{H}_{q}$–module, and every simple $\mathcal{H}_{q}$–module arises this way [3, II, §11D]. On modules and characters, the correspondence is
@@
This identification extends $\mathbb{Z}$–linearly and preserves parabolic induction [6, §9.1.9]: $1_{P_{I}}^{G}=\mathbb{C}Ge_{P_{I}}\leftrightarrow 1_{W_{I}}^{W_{n}}$.
@@
    For $W_{n}$, Lemma 6.1.3 of [6] gives, for $(\lambda_{1},\mu_{1})\vdash n_{1}$, $(\lambda_{2},\mu_{2})\vdash n_{2}$,
@@
    A further useful identity (Lemma 6.1.4 of [6]) is $1_{S_{n}}^{W_{n}}=\sum_{i=0}^{n}\chi^{((i),(n-i))}.$
@@
Now let $x=\sum_{g\in W_{n}}c_{g}g\in Z(\mathbb{C}W_{n})$, and denote by $x^{\prime}=\sum_{g\in H}c_{g}g\in\mathbb{C}H$ its restriction. By Schur’s lemma, $x$ acts on $S^{\lambda,\mu}$ by some scalar $c$, while $x^{\prime}$ acts on $\widetilde{S}^{\lambda,\mu}$ by a scalar $c^{\prime}$. Since $S^{\lambda,\mu}=\mathbb{C}W_{n}\otimes_{\mathbb{C}H}\widetilde{S}^{\lambda,\mu}$, for any $m\in\widetilde{S}^{\lambda,\mu}$ we have
@@
But $1\otimes(x^{\prime}m)=c^{\prime}\,(1\otimes m)$, hence $c=c^{\prime}$. Therefore $x$ acts on $S^{\lambda,\mu}$ by the same scalar as $x^{\prime}$ acts on $\widetilde{S}^{\lambda,\mu}$.
@@
    For bipartitions $(\lambda,\mu)\vdash n$ these polynomials are given by the hook–cohook formula [2, §13.8]. To describe it, one forms the defect–1 symbol $S=(X,Y)$ of $(\lambda,\mu)$ ([2, §11.4.2]):
@@
    Then the formula in [2, §13.8] simplifies to
@@
    By the induction–restriction scalar lemma above, $z_{\mathrm{long}}$ acts on $S^{\lambda,\mu}$ by the same scalar as $z^{\prime}_{\mathrm{long}}$ acts on $\widetilde{S}^{\lambda,\mu}$. Since $\rho_{n_{+},n_{-}}(1+\tau_{i}\tau_{j})=2$ for $i,j$ within the same block, it follows that
@@
    Block factorization Let $|\mu|=n_{-}=2j$ (if $n_{-}$ is odd the character value vanishes) and $|\lambda|=n-n_{-}$. Group the minus coordinates into two blocks $J_{3},J_{4}$ of size $j$, and the plus coordinates into two blocks $J_{1},J_{2}$ of size iji-j together with $K_{1}$ of size $k$. Let $u_{2}$ swap $J_{3}\leftrightarrow J_{4}$ and $u_{1}$ swap $J_{1}\leftrightarrow J_{2}$ while fixing $K_{1}$ pointwise. Denote by $e_{J_{r}}$ the averaging idempotent of $S_{J_{r}}$. Then a direct, but somewhat lengthy, theoretical computation shows that
@@
    Values. Write $\mu=(\mu_{1},\mu_{2})$ and $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$ in descending order. For the $\mu$ part, it is known [18, Prop. 8.5.2] that $\chi^{\mu}(u_{2}e_{J_{3}}e_{J_{4}})=(-1)^{\mu_{1}}$. For the $\lambda$ part, write $|\lambda|=2\ell+k$, then

### v3

@@ 3.4 Luo’s decomposition theorem
The following subsection concerns a recent result of Luo [undefo, §2] on the decomposition of the longest element $w_{S}$ of a Weyl group. Luo originally formulated the theorem in the context of his work on the Casselman–Shahidi conjecture, where it serves as a key combinatorial input for analyzing the singularity of intertwining operators in exceptional types. In our setting the same theorem provides the structural simplification needed to collapse long products of relative destination elements to just a few double coset operators involving opposite parabolics.
@@
In terms of our notations, Luo’s original statement ([undefo]) is
@@
with length additivity. This is equivalent to $w_{1}w_{S}=(w^{\prime}_{1}w_{2})\cdots(w^{\prime}_{m-1}w_{m})$, which is in turn equivalent (by multiplying $(w^{\prime}_{0}w_{1})$ on the left) to the form
@@
used above. We adopt the latter form because it aligns more directly with the orbit structure of the geodesic edge graph and with our RDE computations, and because $2m$ is a period for the evolution of types. The values of $m$ for each irreducible type of $\Delta$ are listed in next section.
@@ 4 Examples: Explicit Zeta Factors for Finite Buildings
This section illustrates the general spectral and zeta–factor formulas developed in Section 3.2 by working out concrete cases for spherical buildings of finite classical groups. We focus on the split families of type $\mathbf{A}$ and type $\mathbf{C}$, for which all Hecke parameters equal $q$ and the zeta factors admit closed expressions. The type $\mathbf{A}$ case recalls the results of [undefr] in the present framework, while the type $\mathbf{C}$ case provides new explicit formulas for $G=\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$.
@@ 4.1 Finite buildings: split and non–split cases
By Tits’ classification of buildings, every finite spherical building of irreducible type and rank at least three arises from an algebraic group of Lie type over a finite field [undefu, §11.4]. A complete list of such families is given in [undefu, §11.2]. Our discussion applies uniformly to all of them.
@@
For twisted or non–split groups, the parameters $Q(s)$ may vary across different conjugacy classes of simple reflections. An important infinite family is provided by the unitary groups $^{2}A_{n}(q)=\mathrm{PGU}_{n+1}(q)$ ($n\geq 2$), defined by a non–degenerate hermitian form. Their building is the flag complex of the associated polar space, of type $\mathbf{C}_{\lfloor(n+1)/2\rfloor}$ [undefu, §11.2.2].
@@
$Z_{c}\!\bigl(X_{2}(B),u\bigr)=\prod_{C}Z_{C}(u),\qquad Z_{C}(u)=\prod_{\lambda\in\mathop{\rm Spec}\nolimits(T_{C})}\frac{1}{1-\lambda\,u^{\,c}}.$
@@
$Z_{c}\!\left(X_{2}(\mathcal{B})\!\mid_{C},u\right)=\prod_{\chi:\,n_{\chi}\neq 0}\;\prod_{\zeta^{d}=1}\frac{1}{\bigl(1-\zeta\,(Q(w_{I})^{-2}q^{\,f_{\chi}})^{1/d}\,u^{c}\bigr)^{m_{C,\chi}(\zeta)\,d_{\chi}}}.$
@@ 4.2 Type $\mathbf{A}$: recalling the formulas
We summarize the zeta functions for buildings of type $\mathbf{A}_{n}$ from [undefr] using the present three–step method; we refer to that paper for the details. By Tits’ classification [undefu, §11.2], all buildings of type $\mathbf{A}_{n}$ with $n\geq 3$ are associated with $G=\mathop{\rm GL}\nolimits_{n+1}(\mathbb{F}_{q})$.
@@
Step 3. Root selection. The eigenvalues of $T_{C}$ are the $(2m/c)$-th roots of the collapsed values above. In type $A$, $2m=6$ (See for example §4.4 below).
@@
  • If $c=2$ (balanced), then $2m/c=3$: we need to take cubic roots of the eigenvalues $q^{\,f_{\lambda}-2\,\mathrm{wt}_{r}(\mu)}$. Multiplicities are as in Table 8 of [undefr, §8.3]: the $K_{\lambda,\mu}$ eigenvalues split as evenly as possible among $1,\omega,\omega^{2}$ ($\omega^{3}=1$). Thus
@@
    and $m(1)$ differs from $m(\omega)$ by at most $1$, with $m(1)-m(\omega)\equiv K_{\lambda,\mu}\bmod 3$.
Therefore we recover the following closed formulas in type $A$ [undefr, §7.2, §8.4].
@@
$Z_{c}\bigl(X_{2}(B)|_{C},u\bigr)\;=\;\prod_{\lambda\vdash n}\frac{1}{\bigl(1-q^{\,f_{\lambda}-2\,\mathrm{wt}_{r}(\mu)}\,u^{6}\bigr)^{d_{\lambda}K_{\lambda,\mu}}}.$
@@
$\frac{1}{Z_{c}\!\bigl(X_{2}(B)|_{C},u\bigr)}\;=\;\prod_{\lambda\vdash n}\;\prod_{j=0}^{2}\Bigl(1-\omega^{j}\,q^{(f_{\lambda}-2\,\mathrm{wt}_{r}(\mu))/3}\,u^{2}\Bigr)^{d_{\lambda}\,(K_{\lambda,\mu}+\epsilon_{\lambda,j})/3},$
@@ 4.3.1 Geometry and zeta factors set-up
$Z_{c}\!\bigl(X_{2}(\mathcal{B})\!\mid_{C},u\bigr)=\prod_{\chi:\,n_{\chi}\neq 0}\ \prod_{\zeta^{\,d}=1}\frac{1}{\Bigl(1-\zeta\,(Q(w_{I})^{-2}q^{\,f_{\chi}})^{1/d}\,u^{\,c}\Bigr)^{m_{C,\chi}(\zeta)\,d_{\chi}}}.$
@@
Beyond the split symplectic case $G=\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$, the same spherical buildings of type $\mathbf{C}$ arise from, among others, two other classical families [undefu, §11.2].
@@
Along a fixed type–orbit $C$ in the geodesic edge graph, the invariants $\ell(w_{I}),\,f_{\chi},\,n_{\chi},$ and $m_{C,\chi}(\zeta)$ depend only on the Weyl group $W(C_{n})$ and on the chosen orbit $C$; only the degree $d_{\chi}$ depends on the ambient group representation. For the natural correspondence between labels $(\lambda,\mu)\vdash n$ on $W(C_{n})$ and the representations of $G=\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$ or $\mathrm{O}_{2n+1}(\mathbb{F}_{q})$, the associated representations for $\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q})$ and $\mathrm{O}_{2n+1}(\mathbb{F}_{q})$ have the same dimensions $d_{\chi}$ [undefa, §13.8]. Consequently, the edge–geodesic zeta factors for $\mathcal{B}(\mathop{\rm Sp}\nolimits_{2n}(\mathbb{F}_{q}))$ and $\mathcal{B}(\mathrm{O}_{2n+1}(\mathbb{F}_{q}))$ coincide, although the two buildings are not isomorphic when $n\geq 3$.
@@ 4.3.2 Spectral inputs for type $\mathbf{C}$
Let $S^{\lambda},S^{\mu}$ denote the Specht modules of the symmetric group $S_{n_{+}},S_{n_{-}}$, so their characters are $\chi_{\lambda,1}$ and $\chi_{\mu,1}$ in §4.2. Inflate $S^{\lambda}\boxtimes S^{\mu}$ to $C_{2}^{\,n}\rtimes(S_{n_{+}}\times S_{n_{-}})$ (trivial on $C_{2}^{\,n}$) and set
@@
    $\mathop{\rm Irr}\nolimits(\mathcal{H}_{q})$ vs. $\mathop{\rm Irr}\nolimits(W_{n})$ The family $\{\mathcal{H}_{q}\}_{q\neq 0}$ is a flat deformation of $\mathbb{C}W_{n}$: for $q=1$ one has $\mathcal{H}_{1}=\mathbb{C}W_{n}$. By Tits’ deformation theorem [undefe, §8.1.7], there is a correspondence between irreducible representations
@@
    If $M$ is a simple $\mathbb{C}G$–module contained in $\mathbb{C}Ge_{B}$, then $e_{B}M$ is a simple $\mathcal{H}_{q}$–module, and every simple $\mathcal{H}_{q}$–module arises this way [undefb, II, §11D]. On modules and characters, the correspondence is
@@
This identification extends $\mathbb{Z}$–linearly and preserves parabolic induction [undefe, §9.1.9]: $1_{P_{I}}^{G}=\mathbb{C}Ge_{P_{I}}\leftrightarrow 1_{W_{I}}^{W_{n}}$.
@@
    For $W_{n}$, Lemma 6.1.3 of [undefe] gives, for $(\lambda_{1},\mu_{1})\vdash n_{1}$, $(\lambda_{2},\mu_{2})\vdash n_{2}$,
@@
    A further useful identity (Lemma 6.1.4 of [undefe]) is $1_{S_{n}}^{W_{n}}=\sum_{i=0}^{n}\chi^{((i),(n-i))}.$
@@
Now let $x=\sum_{g\in W_{n}}c_{g}g\in Z(\mathbb{C}W_{n})$, and denote by $x^{\prime}=\sum_{g\in H}c_{g}g\in\mathbb{C}H$ its restriction (projection) to $\mathbb{C}H$. Then $x^{\prime}\in Z(\mathbb{C}H)$, and by Schur’s lemma, $x$ acts on $S^{\lambda,\mu}$ by some scalar $c$, while $x^{\prime}$ acts on $\widetilde{S}^{\lambda,\mu}$ by a scalar $c^{\prime}$. Since $S^{\lambda,\mu}=\mathbb{C}W_{n}\otimes_{\mathbb{C}H}\widetilde{S}^{\lambda,\mu}$, for any $m\in\widetilde{S}^{\lambda,\mu}$ we have
@@
But $1\otimes(x^{\prime}m)=c^{\prime}\,(1\otimes m)$, hence $c=c^{\prime}$. Therefore $x$ acts on $S^{\lambda,\mu}$ by the same scalar as $x^{\prime}$ acts on $\widetilde{S}^{\lambda,\mu}$. We will use this scalar equality in (b) and (c) below.
@@
    For bipartitions $(\lambda,\mu)\vdash n$ these polynomials are given by the hook–cohook formula [undefa, §13.8]. To describe it, one forms the defect–1 symbol $S=(X,Y)$ of $(\lambda,\mu)$ ([undefa, §11.4.2]):
@@
    Then the formula in [undefa, §13.8] simplifies to
@@
    By the induction–restriction scalar equality above, $z_{\mathrm{long}}$ acts on $S^{\lambda,\mu}$ by the same scalar as $z^{\prime}_{\mathrm{long}}$ acts on $\widetilde{S}^{\lambda,\mu}$. Since $\rho_{n_{+},n_{-}}(1+\tau_{i}\tau_{j})=2$ for $i,j$ within the same block, it follows that
@@
    Block factorization Let $|\mu|=n_{-}=2j$ (if $n_{-}$ is odd the character value vanishes) and $|\lambda|=n-n_{-}$. Group the minus coordinates into two blocks $J_{3},J_{4}$ of size $j$, and the plus coordinates into two blocks $J_{1},J_{2}$ of size iji-j together with $K_{1}$ of size $k$. Let $u_{2}$ swap $J_{3}\leftrightarrow J_{4}$ and $u_{1}$ swap $J_{1}\leftrightarrow J_{2}$ while fixing $K_{1}$ pointwise. Denote by $e_{J_{r}}$ the averaging idempotent of $S_{J_{r}}$. Then a direct, but somewhat lengthy, theoretical computation shows that (a detailed derivation and generalization will appear in a future work)
@@
    Values. Write $\mu=(\mu_{1},\mu_{2})$ and $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$ in descending order. For the $\mu$ part, it is known [undefr, Prop. 8.5.2] that $\chi^{\mu}(u_{2}e_{J_{3}}e_{J_{4}})=(-1)^{\mu_{1}}$. For the $\lambda$ part, write $|\lambda|=2\ell+k$, then